how do u solve this equation ok so here is the sentance complete the solution of t he eqation find the value of y when x=28
x+3y=28 please help me solve this problem!

Answer by stanbon(57967)

find the value of y when x=28
x+3y=28 please help me solve this problem!
---------------
Substitute for "x" and solve for "y":
-----
x+3y = 28
28 + 3y = 28
3y = 0
y = 0
============
Cheers,
Stan H.